the digits of a 2 digit number differ by5.the number obtained by reversing the digits is 9 less than the thrice the original number.find the number.
 

Grade:

2 Answers

Ahalya
9 years ago
let the original number be ‘‘x*10+y’’ and the number obtained by reversing the digits be ‘’y*10+x’’(ie)’’10x+y and 10y+x’’.
according to the question equation formed is:
10x-y=5  and 10y+x=3(10x+y)-9 (ie) -29x+7y=-9.
As per elimination method
-29x +7y = -9
 70x  -7y = 35
41x=26
x=26/41=0.6341....
 
and substituting the value of x in any of the equation we can get the value of y.
 
10x-y=5
10*26/41-y=5
 
y=(260/41)-5
y=(260-205)/41
y=55/41=141463......
Sunil Raikwar
9 years ago
Hello student, please check the solution of your question given below

let the no. be 10x+y
according to question
y-x=5.........................1
and
3(10x+y)=10y+x+9
29x-7y=9...............2
by solving equation 1 & 2 we get
x=2, y=7
so the required no. is 27